Imagine the following open economy: C = 50 + 0.75YD , I = 50 - 0.5i, t = 0.1, G = 100, Im = 0.1Y, Ex = 100. Calculate the slope of the IS curve.
1
Expert's answer
2016-10-14T14:45:03-0400
Y = C + I + G + NX, YD = Y-tY, NX = Ex – Im =>

=> Y = 50 + 0.75*(Y – 0.1Y) + 50 – 0,5i + 100 + 100 – 0.1Y =>

=> Y – 0.75Y + 0.075Y + 0.1Y = 50 + 50 – 0.5i + 100 + 100 =>

=> 0.425Y=300 – 0,5i => Y = 705.9 – 1.18i, -1.18 – the slope of IS curve